OpenStudy (anonymous):

The difference of 3 times 4 and 2

OpenStudy (anonymous):

3(4)-2=10

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

well, the equation would be 4 times 3 minus two and in equations you multiply before you add or subtract so yes i am sure :) 3(4)-2=10 then you would do 12-2=10

OpenStudy (anonymous):

its not pemdas

OpenStudy (anonymous):

in order to find this problem you should be using an equation. In equations you do us ethat system of problem solving if im not mistaken and im pretty sure im right.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome :)